# Conditional Generative Modeling for Enhanced Credit Risk Management in Supply Chain Finance

### v1

@@ Abstract
The rapid expansion of cross-border e-commerce (CBEC) has created significant opportunities for small and medium-sized enterprises (SMEs), yet financing remains a critical challenge due to SMEs’ limited credit histories. Third-party logistics (3PL)-led supply chain finance (SCF) has emerged as a promising solution, leveraging in-transit inventory as collateral. We propose an advanced credit risk management framework tailored for 3PL-led SCF, addressing the dual challenges of credit risk assessment and loan size determination. Specifically, we leverage conditional generative modeling of sales distributions through Quantile-Regression-based Generative Metamodeling (QRGMM) as the foundation for risk estimation. We propose a unified framework that enables flexible estimation of multiple risk measures while introducing a functional risk measure formulation that systematically captures the relationship between these risk measures and varying loan levels, supported by theoretical guarantees. To capture complex covariate interactions in e-commerce sales data, we integrate QRGMM with Deep Factorization Machines (DeepFM). Extensive experiments on synthetic and real-world data validate the efficacy of our model for credit risk assessment and loan size determination. This study represents a pioneering application of generative AI in CBEC SCF risk management, offering a solid foundation for enhanced credit practices and improved SME access to capital.
@@ 1 Introduction
E-commerce has become a cornerstone of global trade, with digital platforms enabling businesses of all sizes to reach international customers. In recent years, online retail has seen explosive growth worldwide: Business-to-Consumer (B2C) e-commerce sales have increased from approximately \$1.5 trillion in 2014 to over \$6 trillion in 2024, exceeding the GDPs of Japan and Germany (Huria 2019, World Bank 2025). Within this boom, cross-border e-commerce (CBEC) has emerged as a rapidly growing segment, increasing from around \$300 billion in 2015 to over \$1.2 trillion in 2024 (Future 2025). Despite this scale, access to financing remains a significant barrier—especially for small and medium-sized enterprises (SMEs). For instance, although over 70% of Amazon’s merchants source products internationally, only 1% obtain financial support from traditional banks (Jungle Scout 2024). These financing gaps are driven by SMEs’ limited collateral, thin credit histories, and the complexities of cross-border operations (Chiu 2024, Cai 2024).
@@
Supply chain finance (SCF), particularly the third-party logistics (3PL)-led model, has emerged as a promising solution. In this model, 3PL providers collaborate with financial institutions to manage both logistics and financing, offering inventory-backed loans by leveraging real-time in-transit goods as collateral (Chen and Cai 2011, Ji et al. 2023). A salient example is UPS Capital, which combines logistics infrastructure with financial services (Huang et al. 2019).
@@
Despite the promise of 3PL-led SCF, the practice of CBEC-oriented supply chain finance remains in its early stage, with fragmented adoption and limited standardization across markets. Industry developments reflect this nascent state: for instance, in 2024 the Hong Kong Export Credit Insurance Corporation just launched the market-first trade credit insurance programs specifically tailored to Hong Kong-based CBEC loans (Chiu 2024).This paper focuses on a specific 3PL-led CBEC-SCF mode where SMEs must pre-stock inventory in overseas warehouses to meet consumer expectations for rapid delivery, resulting in substantial working capital pressures.
@@
In this 3PL-led CBEC-SCF mode, SMEs consign goods to 3PL-managed overseas warehouses, which serve as collateral for financial institutions. Lenders then assess credit risk based on the real-time value of in-transit inventory and predictive analytics of sales performance. The loan amount is adaptively adjusted: if sales forecasts indicate high volatility or low market demand, lenders may set loan-to-value ratios as low as 20% or reject financing entirely, while favorable forecasts may lead to higher loan amounts. This loan is secured against the products, which are managed within the 3PL’s coordinated logistics network. As the products eventually delivered, sales transactions on e-commerce platforms generate real-time cash flows. The revenue generated from these sales is then allocated primarily to repay the loan extended by the 3PL (with any surplus then remitted to the SME). This mode is akin to the CBEC-SCF modes proposed by Shi et al. (2023) and Ji et al. (2023), with the added innovation of lenders utilizing predictive sales analytics for credit risk control.
@@
A central challenge in any lending model (and especially in CBEC-SCF) is effective risk management (Qiao and Zhao 2023). Two risk management questions stand out: (1) Credit risk assessment – how to reliably evaluate the probability of a borrower (SME) defaulting on a loan and estimate the potential financial loss and interested risk measures, and (2) Loan size determination – how to set credit limits or loan amounts in line with the borrower’s repayment capacity. These two questions are inherently interrelated and jointly shape the overall risk profile of lending decisions: the outcome of one directly affects the other. As such, addressing them in an integrated manner is essential for robust risk management.
@@
In the 3PL-led CBEC-SCF model, where loan repayments depend on future sales revenues, accurate credit risk assessment and loan sizing hinge on modeling the conditional sales distribution $P(Y|\mathbf{x})$, given contextual features $\mathbf{x}$ such as product attributes (e.g., ratings, pricing, category), seller characteristics (e.g., review scores, platform tenure, fulfillment efficiency), and market factors (e.g., buyer preferences, seasonality). Traditional methods often provide only point estimates like the mean of sales $Y$ (Hastie et al. 2009, Mohri et al. 2018), whereas learning the full conditional distribution $P(Y|\mathbf{x})$ enables a nuanced view of sales variability, including dispersion and tail risk. This richer representation is key for assessing default probabilities and expected losses, as it captures how input features $\mathbf{x}$ influence the full range of potential sales outcomes $Y$, thereby supporting more informed and robust credit decisions.
@@
We identify two promising and distinctive opportunities within this 3PL-led CBEC SCF mode. The first opportunity arises from the immense volume and richness of sales data from e-commerce platforms. Such data come with a variety of easily collectible features, offering a granular view of consumer behavior and sales trends. In contrast, existing credit risk assessment methods in SCF often center on the “SMEs itself-oriented” factors (e.g., credit histories, financial ratios, and operational metrics) and the “supply chain finance-oriented” factors (e.g., core enterprise credit ratings, industry trends, and supply chain concentration) to predict default risk via various classification models (Zhu et al. 2019, Hou et al. 2024, Belhadi et al. 2021). These conventional approaches face several limitations: the sample size is often limited (data about such diverse range of SCF influencing factors is frequently scattered across multiple entities, especially for SMEs); defaults are highly imbalanced (given the low incidence of bankruptcies), decreasing the accuracy of models (Hou et al. 2024); moreover, most of these models assume a fixed credit loan amount, lacking the flexibility required for adaptive loan sizing.
@@
The second opportunity emerges from transformative advances in generative AI, which enable learning the entire conditional distribution critical for risk quantification. While early machine learning excelled at point prediction—estimating conditional expectations through supervised learning (Hastie et al. 2009, Mohri et al. 2018)—such approaches inherently obscure outcome variability. Generative models fundamentally redefine this paradigm. Pioneered by architectures like Generative Adversarial Networks (GANs, Goodfellow et al. 2014) and its conditional variations (CGAN, Mirza and Osindero 2014) and later advanced through Generative Pre-trained Transformers (GPTs, Radford et al. 2018), these generative models are designed to directly learn the entire conditional distribution of $Y$ given $\mathbf{x}$.
@@
Given these opportunities, there is now a clear need for an accurate, robust, and efficient generative model that can learn the conditional distribution of the sales for specific products and covariates. Despite significant progress in generative AI—as highlighted by Hong et al. (2023)—most existing generative models are primarily designed for high-dimensional data (e.g., images, texts, audios, and videos), with an emphasis on generating outputs that are visually or semantically realistic rather than capturing the full underlying data distribution. In the context of SCF risk management, however, accurately and robustly acquiring the conditional distribution of low-dimensional variables like sales volume is critical. Poor forecasting can lead to critical misjudgments: overestimating the right tail of the distribution may result in a loan amount that exceeds the SME’s actual repayment capacity, heightening default risk, while underestimating sales potential might leave the SME under-financed, curtailing its growth and causing lenders to miss safe lending opportunities.
@@
However, popular methods based on GAN structures, such as the CWGAN with gradient penalty (CWGAN-GP, Athey et al. 2021), often do not exhibit the desired level of accuracy and robustness in this setting (Hong et al. 2023). The current state-of-the-art technology for accurately and robustly capturing low-dimensional conditional distributions is Quantile-Regression-based Generative Metamodeling (QRGMM, Hong et al. 2023). In addition to its precision and robustness, QRGMM provides a significant advantage of efficiently generating a substantial number of random observations in real time, which is particularly valuable for estimating risk measures, facilitating the use of Monte Carlo methods to calculate various risk measures of interest, thereby enhancing our flexibility while minimizing estimation errors caused by sampling variance. In this paper, we explore the application of QRGMM for SCF credit risk management, aiming to leverage its capabilities to overcome the limitations inherent in traditional credit risk assessment methods.
@@
While QRGMM offers a promising tool for learning the conditional distribution of sales volume, its application to SCF credit risk management is far from straightforward and presents several non-trivial challenges. Conventional SCF risk models are largely grounded in binary classification, estimating static default probabilities that fail to account for varying loan sizes or richer risk profiles. Even recent generative approaches using GANs focus on data augmentation, remaining confined to classification frameworks and failing to capture responsive loan size-risk relationships (Zhang et al. 2025). In real-world SCF settings, financial institutions demand adaptive tools that capture how risk metrics evolve with different loan levels, enabling real-time, data-driven loan sizing. Bridging generative modeling with practical SCF credit risk thus requires more than data augmentation: it demands a unified framework that translates the full conditional sales distribution output of QRGMM into actionable, loan-level-specific risk measures.
@@
To meet this challenge, we develop an innovative functional formulation that effectively links the learned sales distribution to various credit risk assessments. Specifically, we introduce a unified framework that systematically formulates key risk metrics as explicit functions of the loan level: $r_{1}(l)$ for default probability, $r_{2}(l)$ for expected loss, and the versatile $r_{3}(l)$ for more generalized risk measures. By doing so, we provide a clear and efficient bridge between the generative sales volume model and a broad class of credit risk measures relevant to SCF. Importantly, this formulation accommodates a wide range of risk metrics within a unified and coherent structure, eliminating the need to construct and analyze cumbersome and inefficient separate static models for each risk measure individually, and thereby significantly enhancing decision-making flexibility. This flexibility allows practitioners to adaptively assess diverse credit risks across different loan sizes in real time, a capability that has been largely absent in prior methodologies.
@@
Nevertheless, ensuring the reliability of this generative approach requires robust theoretical guarantees, which introduce another separate set of non-trivial challenges. While QRGMM ensures convergence in distribution for the estimated sales distribution, this alone is insufficient to guarantee the accurate estimation of more complex credit risk measures. In particular, key risk metrics like $r_{2}(l)$ and $r_{3}(l)$ involve non-linear, non-smooth transformations of the sales distribution, often expressed through complex expectations. Such structures prevent the direct application of standard convergence results from QRGMM, as ensuring accurate risk estimation requires uniform control not just over distribution estimates but also over the induced expectations across all loan levels. We overcome these difficulties by carefully exploiting the structures inherent to the SCF setting. Finally, we establish uniform convergence results for all formulated risk measures building upon the linear quantile regression model of Hong et al. (2023), providing a solid theoretical foundation for our unified framework, ensuring its utility and credibility for SCF credit risk management.
@@
In practice, the above linear assumption seems to be rigid, potentially overlooking the complex relationships between product sales and their covariates in SCF. This complexity arises from the multifaceted nature of sales, which are influenced not only by the products and their attributes but also by various combinations of products, stores, and other factors affecting sales distributions. For example, in a cross-border e-commerce company that we worked with, product selection and sales strategies executed by diverse groups significantly affect outcomes due to varied skills and levels of collaboration. Traditional models, such as those used in recommendation systems like matrix factorization (MF, Lee and Seung 2000) and factorization machines (FM, Rendle 2010), primarily focus on predicting expected values rather than distributions. They are effective for learning combinatorial features but fall short in capturing the full range of the targeted distribution. To address this, we propose a novel approach that integrates QRGMM with deep factorization machines (DeepFM, Guo et al. 2017), enhancing accuracy in predicting target distributions. Extensive numerical experiments with both synthetic and real-world data validate our method’s ability to capture complex covariate interactions and precisely predict the conditional distribution of sales, which facilitates the estimation of pertinent credit risk measures and supports well-informed decision-making processes.
@@
The remainder of the paper is structured as follows. Section 2 presents our functional risk formulation for loan-level determination in SCF, followed by Section 3 which introduces the quantile-regression-based generative model for SCF risk management and details its theoretical properties. Section 4 extends our framework by integrating DeepFM to effectively capture complex covariate interactions. Comprehensive numerical experiments with both synthetic and real-world data are presented in Section 5 to validate our approach. Finally, Section 6 concludes the paper and discusses potential avenues for future research.
2 Functional Risk Formulation for Loan-Level Determination in SCF

### v3

@@ Abstract
The rapid expansion of cross-border e-commerce (CBEC) has created significant opportunities for small- and medium-sized sellers, yet financing remains a critical challenge due to their limited credit histories. Third-party logistics (3PL)-led supply chain finance (SCF) has emerged as a promising solution, leveraging in-transit inventory as collateral. We propose an advanced credit risk management framework tailored for 3PL-led SCF, addressing the dual challenges of credit risk assessment and loan size determination. Specifically, we leverage conditional generative modeling of sales distributions through Quantile-Regression-based Generative Metamodeling (QRGMM) as the foundation for risk measures estimation. We propose a unified framework that enables flexible estimation of multiple risk measures while introducing a functional risk measure formulation that systematically captures the relationship between these risk measures and varying loan levels, supported by theoretical guarantees. To capture complex covariate interactions in e-commerce sales data, we integrate QRGMM with Deep Factorization Machines (DeepFM). Extensive experiments on synthetic and real-world data validate the efficacy of our model for credit risk assessment and loan size determination. This study explores the use of generative models in CBEC SCF risk management, illustrating their potential to strengthen credit assessment and support financing for small- and medium-sized sellers.
@@ 1 Introduction
E-commerce has become an increasingly important component of global trade, with digital platforms enabling businesses of all sizes to reach international customers. Over the past decade, online retail has expanded rapidly: Business-to-Consumer (B2C) e-commerce sales have increased from approximately \$1.5 trillion in 2014 to over \$6 trillion in 2024, exceeding the GDPs of Japan and Germany (WorldBank2019ecommerce, WorldBankEGATE). Within this boom, cross-border e-commerce (CBEC) has emerged as a rapidly growing segment, increasing from around \$300 billion in 2015 to over \$1.2 trillion in 2024 (MarketResearchFuture2025). Despite this scale, access to financing remains a significant barrier, especially for small- and medium-sized sellers. For instance, although over 70% of Amazon’s merchants source products internationally, only 1% obtain financial support from traditional banks (junglescout2024). These financing gaps are driven by small- and medium-sized sellers’ limited collateral, thin credit histories, and the complexities of cross-border operations (Chiu2024, cai2024supply).
@@
Supply chain finance (SCF), particularly the third-party logistics (3PL)-led model, has emerged as a promising solution. In this model, 3PL providers collaborate with financial institutions to manage both logistics and financing, offering inventory-backed loans by leveraging real-time in-transit goods as collateral (chen2011joint, ji2023financial). A salient example is UPS Capital, which combines logistics infrastructure with financial services (huang2019impact).
@@
Despite the promise of 3PL-led SCF, the practice of CBEC-oriented supply chain finance remains in its early stage, with fragmented adoption and limited standardization across markets. Industry developments reflect this nascent state: for instance, in 2024, the Hong Kong Export Credit Insurance Corporation just launched the market-first trade credit insurance programs specifically tailored to Hong Kong-based CBEC loans (Chiu2024). This paper focuses on a specific 3PL-led CBEC-SCF model where small- and medium-sized sellers must pre-stock inventory in overseas warehouses to meet consumer expectations for rapid delivery, resulting in substantial working capital pressure.
@@
In this 3PL-led CBEC-SCF model, small- and medium-sized sellers consign goods to 3PL-managed overseas warehouses, which serve as collateral for financial institutions (FI). Lenders then assess credit risk based on the real-time value of in-transit inventory and predictive analytics of sales performance. The loan amount is adaptively adjusted: if sales forecasts indicate high volatility or low market demand, lenders may set loan-to-value ratios as low as 20% or reject financing entirely, while favorable forecasts may lead to higher loan amounts. This loan is secured against the products, which are managed within the 3PL’s coordinated logistics network. As the products are eventually delivered, sales transactions on e-commerce platforms generate real-time cash flows. The revenue generated from these sales is then allocated primarily to repay the loan (with any surplus then remitted to the seller). This model is akin to the CBEC-SCF models proposed by shi4450732evaluating and ji2023financial, with the added innovation of lenders utilizing predictive sales analytics for credit risk control.
@@
A central challenge in any lending model (and especially in CBEC-SCF) is effective risk management (qiao2023highlight). Two risk management questions stand out: (1) Credit risk assessment — how to reliably evaluate the probability of a borrower defaulting on a loan and estimate the potential financial loss and interested risk measures, and (2) Loan size determination — how to set credit limits or loan amounts in line with the borrower’s repayment capacity. These two questions are inherently interrelated and jointly shape the overall risk profile of lending decisions: the outcome of one directly affects the other. As such, addressing them in an integrated manner is essential for robust risk management.
@@
In the 3PL-led CBEC-SCF model, where loan repayments depend on future sales revenues, accurate credit risk assessment and loan sizing hinge on modeling the conditional sales distribution $P(Y\,|\,\mathbf{x})$, given contextual features $\mathbf{x}$ such as product attributes (e.g., ratings, pricing, category), seller characteristics (e.g., review scores, store tenure, fulfillment efficiency), and market factors (e.g., buyer preferences, industry trends, seasonality). Traditional methods often provide only point estimates like the mean of sales $Y$ (hastie2009elements, mohri2018foundations), whereas learning the full conditional distribution $P(Y\,|\,\mathbf{x})$ enables a nuanced view of sales variability, including dispersion and tail risk. This richer representation is key for assessing risk measures like the default probability and the expected loss, as it captures how input features $\mathbf{x}$ influence the full range of potential sales outcomes $Y$, thereby supporting more informed and robust credit decisions.
@@
We identify two promising and distinctive opportunities within this 3PL-led CBEC SCF model. The first opportunity arises from the immense volume and richness of sales data from e-commerce platforms. Such data comes with a variety of easily collectible features, offering a granular view of consumer behavior and sales trends. In contrast, existing credit risk assessment methods in SCF often center on the “small- and medium-sized-enterprise-oriented” factors (e.g., credit histories, financial ratios, and operational metrics) and the “supply-chain-finance-oriented” factors (e.g., core enterprise credit ratings, industry trends, and supply chain concentration) to predict default risk via various classification models (zhu2019forecasting, hou2024predicting, belhadi2025ensemble). These conventional approaches face several limitations: the sample size is often limited (data about such a diverse range of SCF influencing factors is frequently scattered across multiple entities, especially for small- and medium-sized enterprises); defaults are highly imbalanced (given the low incidence of bankruptcies), decreasing the accuracy of models (hou2024predicting); moreover, most of these models assume a fixed credit loan amount, lacking the flexibility required for adaptive loan sizing.
@@
The second opportunity emerges from transformative advances in generative models, which enable learning the entire conditional distribution that plays a critical role in risk quantification. While early machine learning excelled at point prediction, estimating conditional expectations through supervised learning (hastie2009elements, mohri2018foundations), such approaches inherently obscure outcome variability. Generative models fundamentally redefine this paradigm. Pioneered by architectures like Generative Adversarial Networks (GANs, goodfellow2014generative) and its conditional variations (CGAN, mirza2014conditional) and later advanced through Generative Pre-trained Transformers (GPTs, radford2018improving), these generative models are designed to directly learn the entire conditional distribution of $Y$ given $\mathbf{x}$.
@@
Given these opportunities, there is now a clear need for an accurate, robust, and efficient generative model that can learn the conditional distribution of the sales for specific products and covariates. Despite significant progress in generative models, as highlighted by hong2023learning, most existing generative models are primarily designed for high-dimensional data (e.g., images, texts, audios, and videos), with an emphasis on generating outputs that are visually or semantically realistic rather than capturing the full underlying data distribution. In the context of SCF risk management, however, accurately and robustly acquiring the conditional distribution of low-dimensional variables like sales volume is critical. Poor forecasting can lead to significant misjudgments: overestimating the right tail of the distribution may result in a loan amount that exceeds the seller’s actual repayment capacity, heightening default risk, while underestimating sales potential might leave the seller under-financed, curtailing its growth and causing lenders to miss safe lending opportunities.
@@
However, popular generative models based on GAN structures, such as the CWGAN with gradient penalty (CWGAN-GP, athey2021using), often do not exhibit the desired level of accuracy and robustness in this setting (hong2023learning). The current state-of-the-art generative models for accurately and robustly capturing low-dimensional conditional distributions is Quantile-Regression-based Generative Metamodeling (QRGMM, hong2023learning). In addition to its precision and robustness, QRGMM provides a significant advantage of efficiently generating a substantial number of random observations in real time, which is particularly valuable for estimating risk measures, facilitating the use of Monte Carlo methods to calculate various risk measures of interest, thereby enhancing our flexibility for risk measures estimation while minimizing estimation errors caused by sampling variance. In this paper, we explore the application of QRGMM for SCF credit risk management, aiming to leverage its capabilities to overcome the limitations inherent in traditional credit risk assessment methods. It is important to note that, throughout the paper, we use the term “generative model” in its broader statistical sense, referring to methods that approximate probability distributions and generate random observations, not limited to recent deep-learning-based architectures.
@@
While QRGMM offers a promising tool for learning the conditional distribution of sales volume, its application to SCF credit risk management is far from straightforward and presents several non-trivial challenges. Conventional SCF risk models are largely grounded in binary classification, estimating static default probabilities that fail to account for varying loan sizes or richer risk profiles. Even recent generative approaches using GANs focus on data augmentation for training classification models, remaining confined to classification frameworks and failing to capture responsive loan size-risk relationships (zhang2025credit). In real-world SCF settings, financial institutions demand adaptive tools that capture how risk measures evolve with different loan levels, enabling real-time, data-driven loan sizing. Bridging generative modeling with practical SCF credit risk management thus requires more than data augmentation: it demands a unified framework that translates the full conditional sales distribution output of QRGMM into actionable, loan-level-specific risk measures.
@@
To meet this challenge, we develop a functional formulation as a bridge that effectively links the learned sales distribution to various credit risk measures relevant to SCF. Specifically, we introduce a unified framework that systematically formulates key risk measures as explicit functions of the loan level: $r_{1}(l)$ for default probability, $r_{2}(l)$ for expected loss, and the versatile $r_{3}(l)$ for more generalized risk measures. Importantly, this formulation accommodates a wide range of risk measures within a unified and coherent structure, eliminating the need to construct and analyze cumbersome and inefficient separate static models for each risk measure individually, and thereby significantly enhancing decision-making flexibility. This flexibility allows practitioners to adaptively assess diverse credit risk measures across different loan sizes in real time, a capability that has been largely absent in prior methodologies.
@@
Nevertheless, ensuring the reliability of this generative approach requires robust theoretical guarantees, which introduce another separate set of non-trivial challenges. While QRGMM ensures convergence in distribution for the estimated sales distribution, this alone does not directly guarantee the accurate estimation of many complex credit risk measures. In particular, key risk measures like $r_{2}(l)$ and $r_{3}(l)$ involve non-linear, non-smooth transformations of the sales distribution, often expressed through complex expectations. Such structures prevent the direct application of standard convergence results from QRGMM, as ensuring accurate risk measures estimation requires uniform control not just over distribution estimates but also over the induced expectations across all loan levels. We overcome these difficulties by carefully exploiting the structures inherent to the SCF setting. Finally, building upon the linear quantile regression model of hong2023learning, we establish uniform convergence results for all formulated risk measures, providing a solid theoretical foundation for our unified framework, ensuring its utility and credibility for SCF credit risk management.
@@
In practice, the above linear assumption seems to be rigid, potentially overlooking the complex relationships between product sales and their covariates in SCF. This complexity arises from the multifaceted nature of sales, which are influenced not only by the products and their attributes but also by the combinatorial effects of products, stores, and other contextual factors that shape sales distributions. For example, in a cross-border e-commerce company that we worked with, product selection and sales strategies executed by diverse groups significantly affect outcomes due to varied skills and levels of collaboration. Traditional models, such as those used in recommendation systems like matrix factorization (MF, lee2000algorithms) and factorization machines (FM, rendle2010factorization), primarily focus on predicting expected values rather than distributions. They are effective for learning combinatorial features but fall short in capturing the full range of the targeted distribution. To address this, we propose a novel approach that integrates QRGMM with deep factorization machines (DeepFM, guo2017deepfm), enhancing accuracy in predicting target distributions. Extensive numerical experiments with both synthetic and real-world data validate our method’s ability to capture complex covariate interactions and precisely predict the conditional distribution of sales, which facilitates the estimation of pertinent credit risk measures and supports well-informed decision-making processes.
The main contributions of this paper can be summarized as follows.
  • We develop a generative modeling framework for supply chain finance that provides a distributional characterization of sales and a unified analytical formulation linking the sales distribution to credit risk assessment and loan sizing.
  • We establish uniform convergence results for a broad class of credit risk measures, provide theoretical guarantees that extend beyond distributional convergence, and ensure consistency across the loan decision space.
  • We extend QRGMM with DeepFM to capture complex feature interactions while retaining sampling efficiency, and demonstrate through extensive experiments that it outperforms QRGMM and alternative generative approaches in distributional accuracy, risk measures estimation, and stability.
@@
The remainder of the paper is structured as follows. Section 2 presents the 3PL-led CBEC-SCF model and our functional risk formulation, followed by Section 3, which introduces the quantile-regression-based generative model for SCF risk management, and Section 4 details its theoretical properties. Section 5 extends our framework by integrating DeepFM to effectively capture complex covariate interactions. Comprehensive numerical experiments with both synthetic and real-world data are presented in Section 6 to validate our approach. Finally, Section 7 concludes the paper and discusses potential avenues for future research.
2 3PL-led CBEC SCF Model and Functional Risk Formulation
This section presents both the business model of 3PL-led CBEC-SCF and its functional risk formulation. Section 2.1 details the financing model and illustrates the flow of goods, cash, data, and control among the key actors. Section 2.2 then develops the mathematical risk formulation, linking the operational mechanism to quantitative risk analysis.
2.1 3PL-led CBEC SCF Model
We begin by outlining the 3PL-led CBEC supply chain finance model, which describes the interactions among the seller, the e-commerce platform, the third-party logistics provider, and the financial institution. Specifically, we explain how this model operates by tracing the flows of goods, cash, data, and control, and by highlighting the financial institution’s role in coordinating financing and data-driven risk management. Finally, we contrast this approach with other SCF practices to clarify its distinctive features and positioning.
In practice, the 3PL-led CBEC SCF process unfolds as follows. The seller first pre-positions goods in an overseas warehouse operated by the 3PL provider. These goods are pledged as collateral, and the warehouse provides real-time visibility into inventory movements and stock levels. Customers place orders on the e-commerce platform and pay the platform directly. The 3PL warehouse then fulfills the orders. The platform routes the seller’s proceeds into a repayment account controlled by the financial institution or by a payment service provider (PSP) appointed by the financial institution. From this account, principal and interest are automatically deducted before any surplus is remitted to the seller, which is a standard “cash-dominion” or lockbox arrangement in asset-based lending. In parallel, the financial institution advances loans to the seller to support working-capital needs, with loan sizes determined on the basis of projected sales and the value of the pledged inventory.
This structure does not require direct information sharing between the 3PL and the e-commerce platform. Instead, the financial institution orchestrates the process: it obtains the seller’s authorization to access the platform’s sales and settlement data through the platform’s official application programming interfaces, e.g., Amazon’s Selling Partner API (AmazonSellingPartnerAPI). Without access to these data, the financial institution cannot properly assess repayment risk and may therefore decline to extend credit to the seller. Beyond that, the financial institution can also build a large database of historical sales records through licensed dataset purchases, strategic partnerships, or experience accumulated from previous lending, which provides a valuable corpus for credit-risk modeling.
This financial-institution-led design of data and cash-flow control sidesteps potential data-sharing frictions between independent logistics providers and platforms. It also accommodates cases in which logistics or platform operators offer their own financing, such as the supply chain finance services of UPS Capital or Amazon Lending. Figure 1 illustrates the corresponding flows of goods, cash, data, and control among the seller, platform, 3PL provider, financial institution, and customers.
Figure 1: Schematic of the 3PL-led CBEC-SCF process.
Compared with other SCF models, the 3PL-led CBEC structure offers several advantages:
  • It enables financing even for sellers with limited credit histories. Unlike buyer-led receivables programs (reverse factoring) that depend on the creditworthiness of large buyers, this model enables financing as long as the seller’s e-commerce sales can be verified and monitored.
  • It allows data-driven and adaptive collateralization. Traditional inventory- or receivables-based lending in domestic trade typically relies on a static borrowing base. By contrast, the CBEC model uses e-commerce sales forecasts and goods in transit or stored in 3PL warehouses as collateral, broadening financing access for smaller sellers.
  • It remains applicable to both independent and captive settings. While some logistics or platform operators provide their own lending, such as UPS Capital or Amazon Lending, the financial-institution-led control of data and cash flows captures the broader range of CBEC transactions where logistics and platforms are organizationally separate.
This model aligns with the broader movement toward data-driven inclusive finance, in which alternative information substitutes for traditional collateral or credit files and expands financing opportunities for small- and medium-sized e-commerce sellers.
2.2 Functional Risk Formulation